A volume of .080L of 2.0x10^-3 M Ba(N03)2 (aq) is added to .020L of 5.0x10^-3 M LixSO4(aq). Will a percipitate form?
Answer #1

Consider reaction BaSO 4 (s) Ba 2+ (aq) + SO 42- (aq)

For above reaction , K sp = [ Ba 2+ ] [ SO 42- ] = 1.1 10 -10

We know that precipitation takes place when Ionic product > solubility product.

To check precipitation , we need to calculate ionic product.

First we need to calculate initial concentration of Ba 2+ & SO 42- ions in the mixed solution before any reactions start.

After mixing two solutions , total volume of solution is 0.080 L + 0.020 L = 0.100 L

We can use dilution formula to find out initial concentrations.

Initial concentration of Ba 2+ = ( 2.0 10 -03 M ) ( 0.080 L / 0.100 L ) = 1.6 10 -03 M

Initial concentration of SO 42- = ( 5.0 10 -03 M ) ( 0.020 L / 0.100 L ) =1.0 10 -03 M

We have Q = [ Ba 2+ ] [ SO 42- ]

Q = ( 1.6 10 -03 ) ( 1.0 10 -03 ) = 1.6 10 -06

Here Ionic product ( 1.6 ​​​​​​​ 10 -06 ) > Solubility product ( 1.0 ​​​​​​​ 10 -10 ) , hence precipitation takes place when we add barium nitrate & lithium sulfate solution.